Session: May 2003 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Sat May 24 13:20:10 2003) Success :

Effelsberg (A. Kraus at Sat May 24 15:49:00 2003) Success :
A few minutes have been lost (used for pointing). After UT1237, all source were below our elevation limit.

Onsala (Michael Lindqvist at Sat May 24 15:49:21 2003) Success :
No known problems. Used new schedule.

Torun (K.Borkowski at Tue May 27 05:02:23 2003) Failure :
Most probably a failure. Some hours before start the PI asked to change the schedule but the operator did not get the .skd file in time, thus he observed with the earlier setup. Until now I did not receive any response from the PI to my query on the fate of thus recorded tape. Probably will be of no use since the frequency was off by about 6MHz.

Jodrell_Bank (Alastair Gunn at Tue May 27 22:10:34 2003) Success :
Used version of schedule provided by PI at the last minute. No known problems.

Westerbork (foley at Mon Jul 7 15:07:48 2003) Minor failures : PHAS
RTB unreliable. Used 13 telescopes in the array

Westerbork (foley at Mon Jul 7 15:15:48 2003) Failure :
OLD schedule run. The version sent the evening before was not seen as the Dwingeloo email system was down